OpenStudy (anonymous):

Evaluate 6x ÷ 3y - 4, if x = 20 and y = 2.

OpenStudy (anonymous):

60

OpenStudy (anonymous):

that doesnt seem to be an answer choice.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

yes . Evaluate 6x ÷ 3y - 4, if x = 20 and y = 2. A. 16 B. 20 C. 36 D. 8

OpenStudy (anonymous):

6(20)/ 3(2) - 4 =

OpenStudy (anonymous):

6 * 20 / 3*2 - 4 = 120 / 6 -4 =

OpenStudy (anonymous):

16

OpenStudy (anonymous):

oh crap

OpenStudy (shaik0124):

a is annswer

OpenStudy (anonymous):

@Coolsector thanks (:

OpenStudy (anonymous):

yw;)